OpenStudy (anonymous):

A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6)?

OpenStudy (anonymous):

@xMissAlyCatx help!!

OpenStudy (alekos):

Start with f(2) =f(1) - 8 = 100 - 8 = 92 and go from there with f(3) etc all the way to f(6)

OpenStudy (anonymous):

is it 52

OpenStudy (alekos):

sorry been away. just one minute

OpenStudy (alekos):

No it's 60